For the ellipse represented by 25x^2 + 16y^2-160y =0 find the center
Answer by stanbon(75887)

For the ellipse represented by 25x^2 + 16y^2-160y =0 find the center
-------
25x^2 + 16(y^2-10y+25) = 16*25
-------
25(x-0)^2 + 16(y-5)^2 = 400
----
Center at (0,5)
===========================
cheers,
Stan H.
============================